OpenStudy (anonymous):

320.17 - (6101.4/1.8)

OpenStudy (anonymous):

Answer in proper significant figures.

OpenStudy (anonymous):

ive tried to answer this four times and missed it all 4 times

OpenStudy (anonymous):

missing something i guess

OpenStudy (anonymous):

yes i understand that

OpenStudy (anonymous):

i know when i multiply I take the smallest amount of sig fig

OpenStudy (anonymous):

and when i add or subtract i take the number of least certainty in common

OpenStudy (anonymous):

its when i do both im getting confused

OpenStudy (anonymous):

i assume i do the parenthesis first

OpenStudy (anonymous):

which comes out to 3389.67

OpenStudy (anonymous):

do i round that to two sig figs or subtract first?

OpenStudy (anonymous):

so the first answer is 3400

OpenStudy (anonymous):

then i take 302.17 - 3400

OpenStudy (anonymous):

-3097.83 or 3098 in sig fig

OpenStudy (anonymous):

well its supposed to be 320.17

OpenStudy (anonymous):

so -3079.83 which would be 3080 in sig fig?

OpenStudy (anonymous):

which i tried that answer and it marked me wrong >.<

OpenStudy (anonymous):

Submitted Answers ANSWER 1: Deduction: -3% 320.17 - (6101.4/1.8) = -3069.50 ANSWER 2: Deduction: -3% 320.17 - (6101.4/1.8) = -3079.83 ANSWER 3: Deduction: -3% 320.17 - (6101.4/1.8) = -3080 ANSWER 4: Deduction: -3% 320.17 - (6101.4/1.8) = -3070

OpenStudy (anonymous):

yes but when you subtract dont you take the last significant decimal place to them both?

OpenStudy (anonymous):

could you show me how you would solve this so i may understand?

OpenStudy (anonymous):

i swear to you im not trying to get anyone to do my homework, but this is frustrating me and i have two more attempts >.<

OpenStudy (anonymous):

yes i understand that

OpenStudy (anonymous):

but here i divide first

OpenStudy (anonymous):

which gives me two sig figs

OpenStudy (anonymous):

33.89.67 = 3400 right?

OpenStudy (anonymous):

3389.67 oops

OpenStudy (anonymous):

then i take 320.17 and subtract 3400

OpenStudy (anonymous):

since 3400 has no decimal spots i round to the nearest 1 spot

OpenStudy (anonymous):

and that how i got 3080

OpenStudy (anonymous):

oh so the two sig figs i get when dividing applies still after i subtract?

OpenStudy (anonymous):

for some reason i thought when i rounded to 3400 i was done with the division part

OpenStudy (anonymous):

but 3400 is ambiguous huh

OpenStudy (anonymous):

still 2 figs

OpenStudy (anonymous):

would it be 3100 or -3100?

OpenStudy (anonymous):

you are my hero, thanks a lot!

OpenStudy (anonymous):

most certainly did

OpenStudy (anonymous):

now i have to do the other three parts of the question but at least I understand a little more

OpenStudy (anonymous):

well i understand mult/div and add/subt.... a little confused when they use both together

OpenStudy (anonymous):

my book sucks at explaining! and thanks for the pdf

OpenStudy (anonymous):

well its a custom made book for my college made by pearson

OpenStudy (anonymous):

Just says Chem 121 College of Southern Nevada

OpenStudy (anonymous):

first solve the bracket and then do subtraction and finnaly the rules of significant.